A sphere has an excess of 2,88 x 1013 electrons. What is the charge on the sphere?


1
Expert's answer
2021-09-08T08:14:04-0400

Since we know that 1 electron has a charge equal to -1.602 X 10-19 C, we use the conversion factor and find the total charge of the electron excess:


"2.88\times 10^{13}\,{electrons}(\dfrac{-1.602\times10^{-19}\,C}{1\,{electron}})
 \implies 2.88\times 10^{13}\,{electrons}=-4.614\times10^{-6}\,C"


In conclusion, 2.88 X 1013 electrons (the excess on the sphere) have a charge equal to -4.614 X 10-6 C.


Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.